Grade 12Analytical Geometry

please please answer as fast as ….........
The ends of latus rectum of a parabola are (6,7),(6,-1). It`s vertex is???/
with solution

Mrunal Sonawane
7 Years agoGrade 12

1 Answer

Arun
7 Years ago
We know that length of Lattice rectum is 4a ( here "a" is the distance between focus and vertex ) so 4a = 8a=2(distance between focus and vertex) And from end points of Lattice rectum foci is centre of these two point So foci is(6,3)So vertex will be (4,3) since distance between focus and vertex is 2 units So equation of parabola is(Y-3)^2=4×2(x-4)(Y-3)^2=8(x-4)
 
Hope it helps